Solve the system by the method of your choice: x=6y-9, -2x + 12y=18
Answer by MathDazed(34)

For this problem, if you simplify the second equation you will actually get the first equation.


-2x + 12y = 18


-2x +12y -12y = 18 - 12y


-2x = -12y + 18


-2x/-2 = -12y/-2 + 18/-2


x = 6y - 9


Since the equations are the same, then both equations have infinitely many solutions because they are actually are the same lines.